Find value of cot(-19pi/3)
 

Grade:11

1 Answers

Harsh Patodia IIT Roorkee
9 years ago
Hi Student

cot(-19pi/3)= -cot (19pi/3)= -cot (6pi + pi/3)=-cot (pi/3) =-1/\sqrt{}3

….….….….….….….….….….….….….….….….….….….….….….….….….….….….…..